# Automate Strategy Finding with LLM in Quant Investment

### v3

@@ A.5 Category-Based Alpha Selection
1:categories $\mathcal{C}=\{C_{1},...,C_{m}\}$, each $C_{i}$ containing a set of Alphas, confidence threshold $X$
2:Initialize selected Alphas $\mathcal{A}=\{\}$
@@
4:     $\mathcal{A}_{i}=\text{SelectBestAlphas}(C_{i})$
@@
[…]
8:         Final $\text{Final Score}=w_{r}\cdot\text{Risk Score}+w_{c}\cdot\text{Confidence Score}$
9:         if Final $\text{Final Score}>X$ then
10:              $\mathcal{A}=\mathcal{A}\cup\{\alpha\}$
@@ A.6 Dynamic Alpha Strategy Construction
6:     $\text{filtered content}\leftarrow\text{LLM}.\text{Filter}(d)$
[…]
8:     for each category $c\in\text{categories}$ do
@@
11:     end for
12:end for
13:/* Phase 2: Multi-Agent Evaluation /
14:Initialize selected alphas $\mathcal{A}\leftarrow\emptyset$
15:for each category $c\in\mathcal{F}$ do
16:     for each alpha $\alpha_{ij}\in\mathcal{F}c$ do
17:         $\theta_{ij}\leftarrow\text{ConfidenceScoreAgent}(\alpha_{ij},\mathcal{M}^{(t)})$triangleright Statistical evaluation
18:         $\rho_{ij}\leftarrow\text{RiskPreferenceAgent}(\alpha_{ij},\mathcal{M}^{(t)})$triangleright Risk assessment
19:         $\text{score}{ij}\leftarrow w_{c}\cdot\theta{ij}+w_{r}\cdot\rho_{ij}$triangleright Combined score
20:         if $\text{score}{ij}>\tau$ then
21:              $\mathcal{A}c\leftarrow\mathcal{A}_{c}\cup\{\alpha{ij}\}$
@@
24:     if $\mathcal{A}c\neq\emptyset$ then
25:         $\alpha_{c}^{*}\leftarrow\arg\max_{\alpha_{ij}\in\mathcal{A}c}\text{score}{ij}$triangleright Select best alpha in category
26:         $\mathcal{A}\leftarrow\mathcal{A}\cup\{\alpha_{c}^{\}}$
@@
29:/* Phase 3: Weight Optimization /
30:Initialize MLP with architecture $\{|\mathcal{A}|,10,1\}$triangleright Input, hidden, output layers
@@
32:$\mathbf{y}{\text{train}}\leftarrow\text{FutureReturns}(\mathcal{S},t{\text{start}}+1,t_{\text{end}}+1)$
33:$\mathbf{w}\leftarrow\text{TrainMLP}(\mathbf{X}{\text{train}},\mathbf{y}{\text{train}})$triangleright Learn optimal weights
@@
35:     $\alpha_{i}^{(t)}\leftarrow\sum_{j=1}^{|\mathcal{A}|}w_{j}\cdot\alpha_{ij}^{(}t)$triangleright Compute composite alpha
@@
37:return $\alpha^{(t)}=\{\alpha_{1}^{(t)},\alpha_{2}^{(t)},\ldots,\alpha_{n}^{(t)}\}$triangleright Final alpha strategy
@@ A.8 Portfolio Construction Methods
Our investment methodology implements a daily portfolio reconstruction using a top-$k$/drop-$n$ selection framework. At each trading session, we rank all securities by their alpha values—quantitative indicators of expected excess returns—and select the top $k$ stocks for portfolio inclusion. This approach targets securities with the strongest signals, potentially exploiting short-term market inefficiencies. We employ an equal-weighting scheme across selected securities, distributing capital homogeneously among the top candidates, which offers diversification benefits and aligns with our view that alpha signals primarily provide directional indications rather than precise return forecasts.
@@
A.9 Case Study Backtest Results
Figure 5: Jan-Jun 2021 Backtest on CSI300
Figure 6: Jan-Jun 2022 Backtest on CSI300
Figure 7: Jan-Jun 2023 Backtest on CSI300
Figure 8: Jan-Jun 2021 Backtest on SP500
Figure 9: Jan-Jun 2022 Backtest on SP500
Figure 10: Jan-Jun 2023 Backtest on SP500

### v4

@@ A.5 Category-Based Alpha Selection
1:Categories $\mathcal{C}=\{C_{1},\ldots,C_{m}\}$, each $C_{i}$ containing a set of alphas; confidence threshold $X$
2:Initialize selected alphas $\mathcal{A}\leftarrow\emptyset$
@@
4:  $\mathcal{A}_{i}\leftarrow$ SelectBestAlphas($C_{i}$)
@@
6:   $\text{risk\_score}\leftarrow$ RiskPreferenceAgent($\alpha$)
7:   $\text{confidence\_score}\leftarrow$ ConfidenceScoreAgent($\alpha$)
8:   $\text{final\_score}\leftarrow w_{r}\cdot\text{risk\_score}+w_{c}\cdot\text{confidence\_score}$
9:   if $\text{final\_score}>X$ then
10:     $\mathcal{A}\leftarrow\mathcal{A}\cup\{\alpha\}$
@@ A.6 Dynamic Alpha Strategy Construction
6:  filtered_content $\leftarrow$ LLM.Filter($d$)
7:  categories $\leftarrow$ LLM.Categorize(filtered_content)
8:  for each category $c\in$ categories do
9:   seed_alphas $\leftarrow$ LLM.GenerateAlphas($c$)
10:   $\mathcal{F}_{c}\leftarrow\mathcal{F}_{c}\cup$ seed_alphas
11:  end for
12:end for
13:/* Phase 2: Multi-Agent Evaluation */
14:Initialize selected alphas $\mathcal{A}\leftarrow\emptyset$
15:for each category $c\in\mathcal{F}$ do
16:  for each alpha $\alpha_{ij}\in\mathcal{F}_{c}$ do
@@
19:   $\text{score}_{ij}\leftarrow w_{c}\cdot\theta_{ij}+w_{r}\cdot\rho_{ij}$ $\triangleright$ Combined score
20:   if $\text{score}_{ij}>\tau$ then
21:     $\mathcal{A}_{c}\leftarrow\mathcal{A}_{c}\cup\{\alpha_{ij}\}$
@@
24:  if $\mathcal{A}_{c}\neq\emptyset$ then
25:   $\alpha_{c}^{*}\leftarrow\arg\max\limits_{\alpha_{ij}\in\mathcal{A}_{c}}\text{score}_{ij}$ $\triangleright$ Select best alpha in category
26:   $\mathcal{A}\leftarrow\mathcal{A}\cup\{\alpha_{c}^{*}\}$
@@
29:/* Phase 3: Weight Optimization */
30:Initialize MLP with architecture $\{|\mathcal{A}|,10,1\}$ $\triangleright$ Input, hidden, output layers
@@
[…]
@@
35:  $\alpha_{i}^{(t)}\leftarrow\sum_{j=1}^{|\mathcal{A}|}w_{j}\cdot\alpha_{ij}^{(t)}$ $\triangleright$ Compute composite alpha
@@
37:return $\alpha^{(t)}=\{\alpha_{1}^{(t)},\alpha_{2}^{(t)},\ldots,\alpha_{n}^{(t)}\}$ $\triangleright$ Final alpha strategy
@@ A.8 Portfolio Construction Methods
Our investment methodology implements a daily portfolio reconstruction using a top-kk/drop-$n$ selection framework. At each trading session, we rank all securities by their alpha values—quantitative indicators of expected excess returns—and select the top $k$ stocks for portfolio inclusion. This approach targets securities with the strongest signals, potentially exploiting short-term market inefficiencies. We employ an equal-weighting scheme across selected securities, distributing capital homogeneously among the top candidates, which offers diversification benefits and aligns with our view that alpha signals primarily provide directional indications rather than precise return forecasts.